Jack flew from Atlanta to San Diego, against a headwind of 52.5 Km/h, in 4 hours. Two hours later, hew flew back to atlanta, with a tailwind of 52.5 Km/h, in 3.5 hours. Find the speed of the plane in still air.
Answer by rwm(914)

r*t=d
4*(p-52.5)=3.5*(p+52.5)
p=787.5